# Duoduo CLIP: Efficient 3D Understanding with Multi-View Images

### v1

@@ 4.4 Fine grained text-to-shape retrieval
Previous evaluations focused on classification or image-based retrieval and do not assess how well the methods align fine-grained text descriptions with 3D models. To evaluate this aspect, we use the test split of Text2Shape [4], which contains $\sim$1.4k tables and chairs from ShapeNet and $\sim$7.4k human-captioned descriptions ($\sim$5 per shape). We report the shape retrieval rates at 1 and 5.
@@
Retrieval performance. We compare the performance of our model to prior work in Table 6. We also include previous state-of-the-art methods that were trained on the Text2Shape dataset. Overall, compared to other methods trained on the Ensembled dataset, our method demonstrates superior performance. The zero-shot performance alone is impressive, surpassing point cloud methods. Notably, we observed a significant discrepancy between the zero-shot CLIP model reported by Ruan et al. [36] and our findings, where the LAION checkpoint trained by OpenCLIP significantly outperforms the original OpenAI checkpoint. Lastly, our model approaches the performance of the SOTA method trained directly on Text2Shape, indicating that there is still much to explore regarding more fine-grained text alignment in large-scale text-to-3D models.
@@
Figure 5: Retrieval using fine-grained text from the Text2Shape dataset. On the left is the input text description and we show the top 3 retrieval results for Uni3D and Duoduo CLIP with the ground truth shape highlighted in green.
@@
Retrieval examples. We present example retrieval results compared to Uni3D in Figure 5. Generally, for text descriptions emphasizing complex textures, such as the ‘man image’ in the first row, our model successfully retrieves the corresponding ground truth 3D model. However, for descriptions focusing on geometric properties, like ‘round back’ in the second row, point cloud-based methods tend to retrieve shapes that better match the specified criteria.
5 Limitations and Future Work
@@
We introduced Duoduo CLIP, an efficient joint text and 3D representation method that leverages multi-view images and surpasses current point cloud based state-of-the-art methods across almost all quantitative metrics while using significantly fewer resources on both synthetic and real datasets.
However, our method sometimes fails in cases requiring geometric understanding, as seen in the qualitative retrieval results from images or text. This is an inherent advantage of point clouds, as each point contains three-dimensional coordinates. To address this, we could explore additional inputs to the model such as depth or normal images of the objects, which provide more geometric information, especially with recent depth estimation methods becoming better. Another potential way to improve performance is with LLMs, as demonstrated by Tang et al. [42], where representations from ULIP-2 were used by an LLM to boost performance and enable question-answering applications. Using our model might further increase performance, offering the flexibility of requiring only a single or a sparse number of images as input. Banani et al. [1] also point out the relative weakness of CLIP pretraining on both 3D structure and consistency benchmarks compared to newer SSL methods [59, 30], which can potentially replace the current pretraining framework used by most works.
Overall, we believe that our method will provide better resource scaling compared to point cloud based methods as more 3D data becomes available in the future.
@@
Acknowledgements. This work was funded by a CIFAR AI Chair, a NSERC Discovery grant, and a CFI/BCKDF JELF grant.

### v3

@@ 4.3 Fine grained text-to-shape retrieval
Previous evaluations focused on classification or image-based retrieval and do not assess how well the methods align fine-grained text descriptions with 3D models. To evaluate this aspect, we use the test split of Text2Shape (Chen et al., 2019), which contains $\sim$1.4k tables and chairs from ShapeNet and $\sim$7.4k human-captioned descriptions ($\sim$5 per shape). We report the shape retrieval rates at 1 and 5.
@@
Retrieval performance. We compare the performance of our model to prior work in Tab. 4. We also include previous state-of-the-art methods that were trained on the Text2Shape dataset. Overall, compared to other methods trained on the Ensembled dataset, our method demonstrates superior performance. The zero-shot performance alone is impressive, surpassing point cloud methods. Notably, we observed a significant discrepancy between the zero-shot CLIP model reported by Ruan et al. (2024) and our findings, where the LAION checkpoint trained by OpenCLIP significantly outperforms the original OpenAI checkpoint. Lastly, our model approaches the performance of the SOTA method trained directly on Text2Shape, indicating that there is still much to explore regarding more fine-grained text alignment in large-scale text-to-3D models.
@@
Figure 7: Retrieval using fine-grained text from the Text2Shape dataset. On the left is the input text description and we show the top 3 retrieval results for Uni3D and Duoduo CLIP with the ground truth shape highlighted in dotted boxes.
@@
Retrieval examples. We present example retrieval results compared to Uni3D in Fig. 7. Generally, for text descriptions emphasizing complex textures, such as the ‘man image’ in the first row, our model successfully retrieves the corresponding ground truth 3D model. However, for descriptions focusing on geometric properties, like ‘round back’ in the second row, point cloud-based methods tend to retrieve shapes that better match the specified criteria.
5 Conclusion
@@
We introduced Duoduo CLIP, which uses a multi-view image encoder to represent shapes and align the embedding to pretrained text-image models. We demonstrated that contrastive learning with multi-view images is an efficient and effective way to encode 3D shapes. Our experiments show that using multi-view images our model surpasses current point cloud state-of-the-art methods, trained in a similar fashion, on both real and synthetic data while using significantly fewer resources.
Multi-view images have limitations. They are potentially poorer at geometric understanding compared to point clouds. One way to alleviate this is to incorporate additional inputs such as depth or normal images, which provide more geometric cues. Our work is also limited in only exploring CLIP-style contrastive pretraining. Our model can be easily adapted to other contrastive pretraining Goel et al. (2022), Zhai et al. (2023) and SSL methods (Zhou et al., 2022, Oquab et al., 2023). This is a promising direction as recent work (Banani et al., 2024) points out the relative weakness of CLIP pretraining on both 3D structure and consistency benchmarks. Another important direction is to examine how to increase the data available for training. Even with a relatively small model as ours, training too many layers can lead to severe overfitting. This suggests that data augmentation and incorporating more 3D data would lead to improved generalization and performance.
By advocating for a reconsideration of multi-view images for shape representation, we believe our work can convince 3D shape researchers to re-examine when point clouds are useful. Images offer several advantages over point clouds, including better capture of texture and color information, less domain gap with pretrained vision models, and better resource scaling. They also offer greater flexibility for practical use cases in real-world applications such as robotics. Leveraging the texture information captured by multi-view methods like ours can complement the geometric understanding of point cloud techniques. A hybrid approach combining both modalities could leverage their respective strengths and adapt to a variety of use cases in the future.
@@
Acknowledgements. This work was funded by a CIFAR AI Chair, an NSERC Discovery grant, and a CFI/BCKDF JELF grant.
Appendix A Appendix
In this appendix, we provide details of inference speed (A.1.1), camera sampling (A.1.2), as well as ablations of the number of multi-views used during training (A.2.1), MLP fine-tuning (A.2.2), and different initialization of encoders (A.2.3). We also include experiments on real-world scanned datasets (A.3), examples of out-of-distribution retrieval (A.4.1), concept mixing (A.4.2), and details of the attention map visualization (A.5.1).
For the ablation experiments in the appendix unless otherwise mentioned, we follow the same default settings as in the main paper. In this default setting, models are trained on the Ensembled dataset and we randomly sample 1 to 6 multi-views during training, with the last 6 attention layers trainable and +MVA enabled using a batch size of 1600.
A.1 Inference speed and camera sampling
A.1.1 Inference Speed
We show the inference speed between different number of trainable layers and frames for -MVA and +MVA in Tab. 8. We benchmark total run time and time per iteration on the entire Objaverse-LVIS (46k shapes). Note that -MVA with 6 layers is not shown, since the run time is the same with -MVA 12 layers during inference since both models have multi-view attention disabled. There is only a slight increase in inference time when +MVA is enabled, and it remains efficient for processing a large number of shapes.
Table 7: Inference speed on Objaverse-LVIS with batch size 200. Default model is highlighted in gray.
Method Layers Frames Speed (s/it) Total Time (s)
+MVA 6 3 0.14 32.80
+MVA 6 6 0.28 65.52
+MVA 6 12 0.56 130.67
-MVA 12 3 0.14 33.23
-MVA 12 6 0.28 63.90
-MVA 12 12 0.54 126.28
+MVA 12 3 0.15 34.01
+MVA 12 6 0.29 68.38
+MVA 12 12 0.59 136.10
Table 8: Different camera sampling settings. Here the elevation is the angle relative to the horizon, with 0 degrees at the horizon, positive angles looking from upward, and negative angles looking from downward.
Setting Radius Azimuth Elevation
UpperHem [1.5, 2.2] [0, 360] [0, 60]
ExtendedHem [1.5, 2.2] [0, 360] [-30, 60]
FullSphere (Zero123) [1.5, 2.2] [0, 360] [-90, 90]
A.1.2 Camera Sampling
To verify the effectiveness of our camera sampling strategy, we benchmark different camera settings (see Tab. 8). Example renders for each of the different settings as well as the quantitative results can be found in Fig. 8. The results show a clear performance improvement as the elevation range increases, suggesting that capturing the object from more extreme angles also provides more context and information about the shape.
Figure 8: Left. We show example renders from each of the different camera settings. Right. We show the Top 1 accuracies of the different camera settings using our default model +MVA (6 layers).
A.2 Other Ablations
A.2.1 Number of Multi-Views during Training
For our default model we choose to sample 1 to 6 multi-views during training as it strikes a good balance between memory usage and performance. In Tab. 10 we show the Objaverse-LVIS accuracy over a variation of sampled views seen during training with other settings kept the same as the default model. Except for the model that is trained with only a single view ie. [1, 1] which uses -MVA as it does not see any multi-views during training.
It can be seen that as the number of multi-views seen during training increases, so does its performance when evaluating at 12 frames during inference. However, the memory usage also rises quickly as attention has quadratic growth when the the number of tokens increases in the case of +MVA. So we choose to sample [1, 6] views as there is an increase of +0.24 on top1 over training with [1, 4] views. However, increasing further to [1, 8] frames only results in a +0.09 top1. Also the amount of memory [1, 6] also allows us to stay within the memory limits of an A5000 (24 GB) GPU.
Table 9: Number of views sampled during training on Objaverse-LVIS ablation. Evaluated using 12 frames per shape during validation. The memory usage is reported on a single GPU with 400 batch size. Default model is highlighted in gray.
Num Frames Train Top1 Top3 Top5 Mem (GB)
[1, 1] 52.97 74.71 81.44 5.14
[1, 2] 54.47 76.37 82.82 9.12
[1, 4] 55.00 76.93 83.36 15.50
[1, 6] 55.24 77.08 83.43 22.18
[1, 8] 55.33 77.30 83.62 28.92
Table 10: Ablation of also training the MLP layers after attention on Objaverse-LVIS, MVPNet and ScanObjectNN. Evaluated using 12 frames of each shape during validation. Default model is highlighted in gray. Trained with batch size 1400 to fit in memory.
Method Layers O-LVIS MVPNet ScanObjectNN
+MVA 6 55.24 49.16 66.32
+MVA 12 55.32 44.42 64.15
+MVA + MLP 6 55.96 45.46 59.63
+MVA + MLP 12 55.63 32.68 60.20
A.2.2 MLP Fine-Tuning
Only fine-tuning the attention layers of the model not only makes the model more efficient to train, but also helps to preserve the priors of the original CLIP model. To showcase this we also enable the training of the MLP layers corresponding to the attention layers being tuned in Tab. 10. Although we find that Objaverse-LVIS performance improves when more of the network (MLP) is allowed to be tuned. It results in worse performance on the other unseen datasets. Our default model performs the best across the other unseen datasets. Note that although our default model also overfits and performs worse on MVPNet compared to the zero-shot model as seen in Tab. 4, the performance drop-off is smaller compared to the other models.
A.2.3 Different Initialization
To understand the importance of initializing the multi-view shape encoder from the same model as the frozen CLIP encoders we conduct additional experiments by using a different initialization. Namely, we use the same ViT-B/32 architecture, but initialize the shape encoder from the OpenAI (Radford et al., 2021) pretrained model and the frozen CLIP model uses the checkpoint laion2b_s34b_b79k from the open source implementation OpenCLIP (Ilharco et al., 2021). Note that we add an additional linear layer for the embeddings of the OpenAI initialized model. The results can be found in Tab. 11.
From the results, we can see that using a different initialization from the frozen clip model results in a large drop-off in performance for both Objaverse-LVIS and ScanObjectNN. This shows the importance of using the same initialization for both shape and frozen CLIP encoders. As the embedding space outputted by the models are already close and allows for smoother fine-tuning and prevents potential "overfitting" to the frozen target text and image embeddings. However, we expect that this problem will diminish with larger datasets, allowing for the use of even more powerful 2D foundational models for the shape encoder such as the DINO models (Caron et al., 2021, Oquab et al., 2023) that also provide better 3D Awareness (Banani et al., 2024).
Table 11: Results of using different initializations for the shape encoder. Same Init follows the settings of the main paper which uses the OpenCLIP pretrained model for both the shape and frozen CLIP encoders. OpenAI Init uses the pretrained model from OpenAI for the shape encoder instead. We evaluate on Objaverse-LVIS and ScanObjectNN with 12 views. We show results for models with different number of trainable layers.
O-LVIS ScanObjectNN
Method Layers Top1 Top3 Top5 Top1 Top3 Top5
OpenAI Init 6 53.0 75.2 82.1 60.8 78.6 86.6
OpenAI Init 12 53.4 75.7 82.5 56.8 78.9 86.8
Same Init 6 55.2 77.1 83.4 66.3 85.5 90.2
Same Init 12 55.3 77.1 83.5 64.2 85.0 90.8
A.3 Other Datasets
A.3.1 ScanObjectNN
To evaluate on ScanObjectNN (a point cloud dataset), we have to first obtain the multi-view images for each object. For simplicity, we rendered the 12 multi-view images using the point clouds (OBJ_BG) directly as can be seen in Fig. 9. Note that ScanObjectNN objects were obtained from real-world scan datasets that were captured with RGBD cameras. So using the raw frames would be more straightforward and easier to obtain compared to point clouds since it does not require reconstruction. However, we weren’t able to obtain the mappings to the original scenes from the processed ScanObjectNN dataset so opted for this approach instead.
We show the quantitative results on ScanObjectNN in Fig. 9. It can be seen that the +MVA mechanism outperforms -MVA by a large margin in this scenario. And that our default model with 12 frames are able to outperform all previous point cloud methods on both top 1 and top 3 accuracies. This is despite the fact that we used rendered point clouds images as opposed to the originally captured RGB scene data. This demonstrates the robustness and adaptability of our model to handle a wide range of different images.
Figure 9: Example of multi-view image renders of ScanObjectNN.
Table 12: ScanObjectNN Results. Our +MVA model significantly outperforms zero-shot CLIP and -MVA. We also outperforms prior work using point clouds, including Uni3D on Top 1 accuracy.
Method Frames Top 1 Top 3 Top 5
ZS B-32 1 37.2 60.3 71.6
ZS B-32 6 50.9 72.4 81.2
ZS B-32 12 53.9 73.5 81.2
-MVA 1 41.2 62.8 74.4
-MVA 6 53.7 75.9 84.4
-MVA 12 56.8 78.4 85.5
OpenShape PC 52.2 79.7 88.7
Recon++ PC 65.4 84.1 89.7
Uni3D PC 65.3 85.5 $\mathbf{92.7}$
+MVA (6 layers) 1 41.7 64.4 76.4
+MVA (6 layers) 6 61.7 82.7 89.2
+MVA (6 layers) 12 $\mathbf{66.3}$ $\mathbf{85.5}$ 90.2
A.3.2 ScanNet
For ScanNet, we generate the multi-view images using the officially released projections from the 3D segmentations onto the original 2D RGB frames. We then sort the frames based on the projected object mask area, selecting the 12 views with the largest areas. Afterward, we crop each object according to its bounding box, centering it within the frame and filling the regions outside the bounding box with white. We compare against recent point cloud based methods in Tab. 13.
It is evident that image-based models significantly outperform point cloud models. Notably, the best-performing model is the zero-shot B-32 CLIP model, highlighting a persistent issue with dataset size and the tendency of these methods to overfit. While we observe improvements from models trained on both Ensembled and MVImgNet datasets, which include natural images, the gains are still small, suggesting that larger datasets could lead to further performance enhancements.
Table 13: ScanNet results with cropped video frames evaluated over 12 multi-views. ZS is the zero-shot B-32 model. Ens and Ens + MVI is the default model trained with ensembled dataset and ensembled with MVImgNet datasets respectively.
Method Avg. Bed Cab Chair Sofa Tabl Door Wind Bksf Pic Cntr Desk Curt Fridg Bath Showr Toil Sink
OpenShape 45.6 66.7 3.2 75.8 83.5 37.7 49.2 47.5 64.9 48.2 1.9 66.1 70.2 1.8 50.0 57.1 45.2 7.1
TAMM 49.4 66.7 4.8 83.6 84.5 48.9 57.9 48.2 80.5 61.3 1.9 60.6 83.6 7.0 41.4 56.1 48.4 3.6
Uni3D 45.8 58.5 3.7 78.8 83.7 54.9 31.3 39.4 70.1 35.1 1.9 27.3 94.2 13.8 38.7 10.7 88.1 47.6
ZS 69.2 83.1 58.2 65.9 90.3 56.3 56.9 43.5 86.2 67.7 44.2 93.7 47.0 79.0 64.5 62.0 99.4 78.6
Ens 61.3 83.1 50.0 46.1 87.2 40.5 53.4 39.6 97.0 45.9 16.0 89.2 65.7 86.6 33.3 56.0 98.0 55.1
Ens+MVI 63.3 87.2 51.3 56.5 88.5 35.1 63.3 49.5 98.7 48.8 25.0 92.9 55.1 77.8 26.9 54.8 98.9 65.0
A.4 Additional Results
Figure 10: We perform out-of-distribution image-based retrieval on Objaverse LVIS by using manually cropped images from a large indoor scene image sourced from the web to retrieve the top 3 matching objects using our DuoduoCLIP model trained with MVImgNet.
A.4.1 Out of Distribution Retrieval
To further assess our model’s understanding of 3D objects, we conduct image-based retrieval using out-of-distribution images. We source indoor scene images from the web and generate additional images using a large text-to-image model (Esser et al., 2024). Objects are manually cropped from these scenes (1 view each) and processed through our shape encoder to obtain query embeddings. For the retrieval database, we compute embeddings for the Objaverse LVIS split using 12 views per object. The top 3 matches are returned based on the highest cosine similarity between the query embeddings and the database embeddings. The results for the image sourced from the web are shown in Fig. 10, while those for the image generated by the text-to-image model are presented in Fig. 11.
Our model successfully retrieves objects that are semantically related to the corresponding cropped portions of the images, demonstrating a strong understanding of object categories and their semantic significance. While the retrieved objects may not visually match the cropped images exactly, it is important to note that these are in-the-wild images, and identical objects may not exist in the Objaverse LVIS dataset.
Figure 11: We perform out-of-distribution image-based retrieval on Objaverse LVIS by using manually cropped images from a large indoor scene image generated from Stable Diffusion 3 (Esser et al., 2024) to retrieve the top 3 matching objects using our DuoduoCLIP model trained with MVImgNet.
A.4.2 Concept Mixing
Figure 12: Additional examples of context mixing where we blend concepts from two distinct objects by identifying the Objaverse shape with the embedding that maximizes similarity to both objects.
To explore the latent space learned by our model, we perform concept mixing by identifying the object embedding within Objaverse that maximizes the similarity to both of the two objects being combined. Similar to OpenShape (Liu et al., 2023a), to blend shape $a$ with shape $b$, we select the object $\hat{i}$ with normalized shape embedding $f^{S}_{\hat{i}}$ such that $\hat{i}=\arg\max_{i}\left(\min(\langle f^{S}_{i},f^{S}_{a}\rangle,\langle f^{S}_{i},f^{S}_{b}\rangle)\right)$. Note that the shape embedding is extracted by encoding the 12 views of each object. The results are illustrated in Fig. 2 and Fig. 12. These examples demonstrate that our model has learned a meaningful latent space capable of blending geometric and semantic concepts. For instance, it can combine a pumpkin and a carriage into a pumpkin carriage (Fig. 2 right) or a horse and a wagon into a horse pulling a wagon (Fig. 2 left). Interestingly, the model can also combine semantic attributes, such as merging a dog and an engine to create a four-legged mechanical creature (Fig. 12 top right) or blending a bird and a bus to form a plane (Fig. 12 bottom left), reflecting an understanding of the geometric function of wings and the concept of transport.
A.5 Additional Details
A.5.1 Attention Map Visualization
For Fig. 4, we visualize attention maps (Abnar & Zuidema, 2020) by first selecting an area of interest in one of the multi-view images (e.g., the plate in the first example). All 12 multi-view images of the object are passed through the model backbone to extract the attention map from the query and key embeddings in the first MVA layer. This produces an attention map of size $(M\times 49)\times(M\times 49)$, where $M$ is the number of multi-views and $49$ is the number of tokens per image (excluding the CLS token, with image resolution 224 and ViT patch size 32, giving $(224/32)^{2}=49$ patches). To retrieve attention specific to the area of interest (e.g., token 11 (plate) in the first image), we extract the corresponding row in the attention map, resulting in an attention response of size $M\times 49$. This response is overlaid onto the original images, as shown in Fig. 4. These visualizations reveal how the attention response to a token in one image corresponds to regions in other multi-view images, demonstrating the model’s ability to capture geometric relationships across views.